[Brahms-l] Shift report 20050522 00:00-08:00

From: <apache@brahms-web.brahms.bnl.gov>
Date: Sun May 22 2005 - 08:03:11 EDT
Shifter leader was rk
Shifters were 


Peaceful night with plenty of data taken.
1. Had beam from the beginning of shift, dump scheduled at 0633 but had been delayed due to problems with LINAC injector - hence still running.
2. Runs: 14825 - 14835 - full field B, FS@2.3, MRS@90.
3. None observed.

_______________________________________________
Brahms-l mailing list
Brahms-l@lists.bnl.gov
http://lists.bnl.gov/mailman/listinfo/brahms-l
Received on Sun May 22 08:03:43 2005

This archive was generated by hypermail 2.1.8 : Sun May 22 2005 - 08:03:54 EDT